OpenStudy (anonymous):

Add and answer in simplest fractional form. -1/5 + -3/4 =

OpenStudy (unklerhaukus):

-19//20

OpenStudy (stormfire1):

-1/5 = -4/20 -3/4 = -15/20 Add the numerators to get -19/20